dear algebra tutors, this baby has stumped me - the 23% confuses me.
one thousand quarts of milk testing 4% butterfat are to be reduced to 3%. how many quarts of cream testing 23% must be separated from the milk to produce the required? all percents are by volume.
thank you!

Answer by checkley71(8403)

1000*.04-.23x=.03(1000-x)
40-.23x=30-.03x
40-30=.23x-.03x
10=.2x
x=10/.2
x=50 quarts of 23% must be removed.
proof
1000*.04-.23*50=.03(1000-50)
40-11.5=30-1.5
28.5=28.5